# Sparse estimation for the drift of high-dimensional Ornstein–Uhlenbeck processes with i.i.d. paths

### v1

@@ 4.1. Setup
The computation is based on the SLOPE package in R (Larsson et al.,, 2025). Initial drafts of some simulation scripts were generated with assistance from ChatGPT 4.0 and then fully reviewed, tested, and revised by the author. All results reported use code that the author verified.
@@ A.2. Eigenvalue behaviours
where $\boldsymbol{w}(t)=(\boldsymbol{w}_{1}(t),\ldots,\boldsymbol{w}_{N}(t))$ is an $Nd$-dimensional standard Wiener process. It clearly holds that $\mathbb{E}[B(h_{1})B(h_{2})]=\langle h_{1},h_{2}\rangle_{\mathbb{H}}=\int_{0}^{T}h_{1}(s)^{\top}h_{2}(s)\mathrm{d}s$ for any $h_{1},h_{2}\in\mathbb{H}$. We define the set of all smooth cylindrical random variables in the form $F=f(B(h_{1}),\ldots,B(h_{n}))$, where $n\in\mathbb{N}$ is arbitrary and $f$ is of $\mathcal{C}^{\infty}$ such that $f$ and its derivatives are of polynomial growth. The Malliavin derivative $DF$ of $F$ is define as
@@
we separately evaluate the first term (first order Wiener chaos) and second term (second order Wiener chaos) on the right-hand side.
(Step i) First, let us consider a bound for the following conditional probability given the initial values:
$\mathbb{P}\left(\left.\left|2\int_{0}^{T}\left(\boldsymbol{X}_{1:N}^{0}(t)\right)^{\top}\mathbf{K}_{\boldsymbol{v}}\left(\boldsymbol{X}_{1:N}^{1}(t)\right)\mathrm{d}t\right|\geq Nu/2\right|\{\boldsymbol{\xi}_{i}\}\right).$
@@
As $Z^{1}(\boldsymbol{v})$ is a first order Wiener chaos (Eq. (4.6) of Nualart and Nualart,, 2018), it holds that $Z^{-1}(\boldsymbol{v})\in\mathbb{D}^{1,2}$, $L^{-1}Z_{1,\boldsymbol{v}}=-Z_{1,\boldsymbol{v}}$, and
@@
$\sum_{i=1}^{N}\left(\left(\boldsymbol{X}_{i}^{0}(s)\right)^{\top}\boldsymbol{v}\right)^{2}=\boldsymbol{v}^{\top}\exp(t\mathbf{A}_{0})\left(\sum_{i=1}^{N}\boldsymbol{\xi}\boldsymbol{\xi}^{\top}\right)\exp(t\mathbf{A}_{0})^{\top}\boldsymbol{v}\leq\left\|\sum_{i=1}^{N}\boldsymbol{\xi}\boldsymbol{\xi}^{\top}\right\|_{\textnormal{op}}\mathfrak{p}_{0}^{2}e^{2T\mathfrak{a}_{0}}.$
Hence, Theorem 4.1 of Nourdin and Viens, (2009) yields a concentration bound such that for some $c\geq 1$, for all $u\geq 0$,
@@
$\mathbb{P}(E):=\mathbb{P}\left(\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\xi}_{i}\boldsymbol{\xi}_{i}^{\top}\right\|_{\textnormal{op}}\leq c(1+\sqrt{d/N}+d/N)\right)\geq 1-2e^{-N}.$
@@
$=\int_{E\cup E^{c}}\mathbb{P}\left(\left.\left|2\int_{0}^{T}\left(\boldsymbol{X}_{1:N}^{0}(t)\right)^{\top}\mathbf{K}_{\boldsymbol{v}}\left(\boldsymbol{X}_{1:N}^{1}(t)\right)\mathrm{d}t\right|\geq Nu/2\right|\{\boldsymbol{\xi}_{i}\}\right)\mathbb{P}(\mathrm{d}\boldsymbol{\xi}_{1})\times\cdots\times\mathbb{P}(\mathrm{d}\boldsymbol{\xi}_{N})$
@@
Since $Z^{2}(\boldsymbol{v})$ is a second order Wiener chaos (Nualart and Nualart,, 2018), it holds that $Z^{2}(\boldsymbol{v})\in\mathbb{D}^{1,2}$, $L^{-1}Z^{2}(\boldsymbol{v})=-Z^{2}(\boldsymbol{v})/2$, and

### v2

@@ 4.1. Setup
The computation is based on the SLOPE package in R (Larsson et al.,, 2025). Initial drafts of some simulation scripts were generated with assistance from ChatGPT 4.0 and then fully reviewed, tested, and revised by the author. All results reported use code that the author verified. The source code and data are available at https://github.com/snakakita/sparse-ou-with-iid-paths.
@@ A.2. Eigenvalue behaviours
where $\boldsymbol{w}(t)=[\boldsymbol{w}_{1}(t)^{\top}\cdots\boldsymbol{w}_{N}(t)^{\top}]^{\top}$ is an $Nd$-dimensional standard Wiener process. It clearly holds that $\mathbb{E}[B(h_{1})B(h_{2})]=\langle h_{1},h_{2}\rangle_{\mathbb{H}}=\int_{0}^{T}h_{1}(s)^{\top}h_{2}(s)\mathrm{d}s$ for any $h_{1},h_{2}\in\mathbb{H}$. We define the set of all smooth cylindrical random variables in the form $F=f(B(h_{1}),\ldots,B(h_{n}))$, where $n\in\mathbb{N}$ is arbitrary and $f$ is of $\mathcal{C}^{\infty}$ such that $f$ and its derivatives are of polynomial growth. The Malliavin derivative $DF$ of $F$ is define as
@@
we separately evaluate the first term (first-order Wiener chaos) and second term (second-order Wiener chaos) on the right-hand side.
(Step i) First, let us consider a bound for the first-order Wiener chaos. Let us define the event
$E_{1}=\left\{\boldsymbol{w}\in\mathcal{C}([0,T];\mathbb{R}^{Nd}),\boldsymbol{\xi}=[\boldsymbol{\xi}_{1}^{\top}\cdots\boldsymbol{\xi}_{N}^{\top}]^{\top}\in\mathbb{R}^{Nd}\colon\left|2\int_{0}^{T}\left(\boldsymbol{X}_{1:N}^{0}(t)\right)^{\top}\mathbf{K}_{\boldsymbol{v}}\left(\boldsymbol{X}_{1:N}^{1}(t)\right)\mathrm{d}t\right|\geq Nu/2\right\}.$
What we need to obtain is an upper bound on $\int_{\mathbb{R}^{dN}}\int_{\mathcal{C}([0,T];\mathbb{R}^{dN})}\mathbbm{1}_{E_{1}}(\boldsymbol{w},\boldsymbol{\xi})\mathbb{P}(\mathrm{d}\boldsymbol{w})\mathbb{P}(\mathrm{d}\boldsymbol{\xi})$ (recall the independence between $\{\boldsymbol{w}_{i}\}$ and $\{\boldsymbol{\xi}_{i}\}$). We begin by bounding the following conditional probability given the initial values: for fixed $\{\boldsymbol{\xi}_{i}\}\subset\mathbb{R}^{d}$,
$\mathbb{P}\left(\left.\left|2\int_{0}^{T}\left(\boldsymbol{X}_{1:N}^{0}(t)\right)^{\top}\mathbf{K}_{\boldsymbol{v}}\left(\boldsymbol{X}_{1:N}^{1}(t)\right)\mathrm{d}t\right|\geq Nu/2\right|\{\boldsymbol{\xi}_{i}\}\right)\left(=\int_{\mathcal{C}([0,T];\mathbb{R}^{dN})}\mathbbm{1}_{E_{1}}(\boldsymbol{w},\boldsymbol{\xi})\mathbb{P}(\mathrm{d}\boldsymbol{w})\right).$
@@
As $Z^{1}(\boldsymbol{v})$ is a first-order Wiener chaos (Eq. (4.6) of Nualart and Nualart,, 2018), it holds that $Z^{-1}(\boldsymbol{v})\in\mathbb{D}^{1,2}$, $L^{-1}Z_{1,\boldsymbol{v}}=-Z_{1,\boldsymbol{v}}$ ($L$ is the generator of the Ornstein–Uhlenbeck semigroup), and
@@
$\sum_{i=1}^{N}\left(\left(\boldsymbol{X}_{i}^{0}(s)\right)^{\top}\boldsymbol{v}\right)^{2}=\boldsymbol{v}^{\top}\exp(t\mathbf{A}_{0})\left(\sum_{i=1}^{N}\boldsymbol{\xi}_{i}\boldsymbol{\xi}_{i}^{\top}\right)\exp(t\mathbf{A}_{0})^{\top}\boldsymbol{v}\leq\left\|\sum_{i=1}^{N}\boldsymbol{\xi}_{i}\boldsymbol{\xi}_{i}^{\top}\right\|_{\textnormal{op}}\mathfrak{p}_{0}^{2}e^{2T\mathfrak{a}_{0}}.$
Hence, Theorem 4.1 of Nourdin and Viens, (2009) yields a concentration bound such that for some $c\geq 1$, for all $u\geq 0$ and $\{\boldsymbol{\xi}_{i}\}$,
@@
$\mathbb{P}(E_{2}):=\mathbb{P}\left(\left\|\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\xi}_{i}\boldsymbol{\xi}_{i}^{\top}\right\|_{\textnormal{op}}\leq c(1+\sqrt{d/N}+d/N)\right)\geq 1-2e^{-N}.$
@@
$=\int_{E_{2}\cup E_{2}^{c}}\int_{\mathcal{C}([0,T];\mathbb{R}^{Nd})}\mathbbm{1}_{E_{1}}(\boldsymbol{w},\boldsymbol{\xi})\mathbb{P}(\mathrm{d}\boldsymbol{w})\mathbb{P}(\mathrm{d}\boldsymbol{\xi})$
$=\int_{E_{2}\cup E_{2}^{c}}\mathbb{P}\left(\left.\left|2\int_{0}^{T}\left(\boldsymbol{X}_{1:N}^{0}(t)\right)^{\top}\mathbf{K}_{\boldsymbol{v}}\left(\boldsymbol{X}_{1:N}^{1}(t)\right)\mathrm{d}t\right|\geq Nu/2\right|\{\boldsymbol{\xi}_{i}\}\right)\mathbb{P}(\mathrm{d}\boldsymbol{\xi}_{1})\times\cdots\times\mathbb{P}(\mathrm{d}\boldsymbol{\xi}_{N})$
@@
Since $Z^{2}(\boldsymbol{v})$ is a second-order Wiener chaos (Nualart and Nualart,, 2018), it holds that $Z^{2}(\boldsymbol{v})\in\mathbb{D}^{1,2}$, $L^{-1}Z^{2}(\boldsymbol{v})=-Z^{2}(\boldsymbol{v})/2$, and